OpenStudy (anonymous):

Given f(x) x-1/2, solve for f^-1 (5) 2 9 11 15 I said its 11 am I right correct if im wrong what is the slope of the lkne which passes through (4,5) and 0,1) undefined 1 0 4 I said it was 1 use the table x -1 ,3 ,5 y 7 , 3,1 Find the rate of change for the given function from x=-1 to ×=5 Idk how to do this one

OpenStudy (anonymous):

Pixk whatever u want to work on I just want someone to check my work n help me on number 3

OpenStudy (bibby):

is this it?\[f(x) = \frac{ x-1 }{ 2 }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The possible answers are at the bottom

OpenStudy (anonymous):

The answers for number 3 is -6 -1 1 6

OpenStudy (bibby):

Start by getting the inverse. call f(x), y. replace y and x and solve for y. \[f(x) = \frac{ x-1 }{ 2 }\] \[y = \frac{ x-1 }{ 2 } -> x = \frac{ y-1 }{ 2 } \]

OpenStudy (triciaal):

f(x) = x^(-1/2) or (x^-1)/2

OpenStudy (anonymous):

The 2nd one u put

OpenStudy (anonymous):

Tricaal

OpenStudy (triciaal):

ok

OpenStudy (bibby):

\[x =\frac{y-1}{2}\] multiply both sides by 2 \[2x ={y-1}\]

OpenStudy (triciaal):

(4,5) and (0,1) slope change in y / corresponding change in x (1-5)/(0-4) =-4/-4 =1